HUBBLE SPACE TELESCOPE DAILY REPORT #1839 PERIOD COVERED: 0000Z (UTC) 03/20/97 - 0000Z (UTC) 03/21/97 Daily Status Report as of 080/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7016 (WF/PC-2 SM97 UV Monitor And Cool Down = Procedure) This operation was used as a post-SM97 cooldown of the WF/PC-2 CCD = to -88C with frequent monitoring of contamination via standard star = GRW+70D582 observations in filter F170W. An internal flat and a dark = calibration image were also taken. The observations were completed as = planed, and no problems were reported 1.2 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1:High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy QDOTULIR17. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.3 Completed NIC1/2/3 7135 (NICMOS Intermediate Focus / Alignment) The NICMOS was used to make observations of NGC-3603. The purpose = of this activity is to perform a series of iterative adjustments of the = NICMOS pupil alignment mechanism to establish the initial optical = alignment in focus and tilt for all three cameras. A dark calibration = image was also taken. The observations were completed as schedule, and = no anomalies were reported. 1.4 Completed Two Sets of FGS 7026 (Fold Flat 3 (FF3) Position = Optimization for FGS-1R) The FGSs were used to make FF3 C FGS-1R S-curves which will be = optimized to achieve a modulation within 10 per-cent of the maximum = accepted value at a particular field location and to produce an acceptable= set of S-curves over the entire FGS field of view. The folding flat was = adjusted using PSTOLs from the ground. The observations were completed = as planned and no problems were reported. 1.5 Completed WF/PC-2 6903 (WF/PC-2 Cycle 6 Calibration: Decontaminatio= n) The WF/PC-2 was used to make several calibration bias = observations, internal flats, K-Spot images and dark calibration images = which are all part of a Cycle 6 decontamination procedure. No science = was interleaved with the dark exposures. The observations were completed = as scheduled, and no problems were reported. 1.6 Completed NIC 1/2/3 7039 (NICMOS to FGS Astrometric Calibration - = Aperture Locations) The NICMOS 1/2/3 cameras were used to make astrometric = observations of the cluster NGC-188 in order to determine the various = aperture locations. The position of NICMOS apertures in the HST focal = plane was measured to an accuracy of +/- 2 arcsec in the V2-V3 plane. = The test were repeated three times to monitor the temporal evolution of = the aperture locations as function of focus position. The observations = were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of STIS/CCD 7092 (CCD Dark Rate and Read Noise) Observations were taken to measure CCD dark current and gather = data for making super darks. The observations were completed as planned, = and no anomalies were reported. 1.8 Completed STIS/CCD 7099 (CCD Flat-Field Stability and Cosmetic = Defect Fraction) These observations were used to verify cosmetic defect fraction on = the CCD using flat-field images, and to investigate flat-field stability = over short (minutes) and long (days)time intervals. The observations were executed as scheduled, = and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: No significant opertional problems were reported. One hundred NICMOS images were dumped from buffer memory by = real-time command at 079/1257Z. Later, at 079/2241Z, the NICMOS filter = wheel movement was enabled. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV activities. /DMH